# Vacancy-Induced Quantum Properties in 2D Silicon Carbide: Atomistic insights from semi-local and hybrid DFT calculations

### v1

@@ III.2 Electronic Properties
HSE | 3.39 | 4.79 | 2.47 | 5.03 |
The electronic band structure of pristine monolayer silicon carbide (1L-SiC) provides fundamental insights into its electronic properties and potential applications. Figure 2 presents a comprehensive comparison of the band structures calculated using four different DFT-XC approaches: PBE, SCAN, r2SCAN, and HSE06. All four functionals consistently predict that 1L-SiC is an indirect bandgap semiconductor, with the valence band maximum (VBM) located at the K-point and the conduction band minimum (CBM) at the M-point of the Brillouin zone. The calculated HSE06 bandgap of 3.39 eV is lower than the previous GW-GGA predictions (3.90 eV) of monolayer SiC [53] and in very good agreement with the HSE06 values of 3.35 eV from [54] and 3.58 eV [55]. The indirect nature of the bandgap is a key feature of 1L-SiC, distinguishing it from other 2D materials like graphene or monolayer transition metal dichalcogenides, and has important implications for its optical and electronic properties. The indirect bandgap in 1L-SiC necessitates phonon assistance for optical transitions to conserve momentum. This three-particle process (electron, hole, and phonon) is less probable than the two-particle process in direct bandgap materials, leading to longer radiative lifetimes. The phonon involvement allows the system to remain in the excited state for a longer duration, as the electron must wait for an appropriate phonon to assist in the recombination process. This mechanism can be advantageous for applications requiring sustained excited states, such as in certain photocatalytic processes or in devices exploiting long-lived excitons. Notably, the spin-up (pink) and spin-down (blue) bands are perfectly degenerate for the pristine case across all functionals, indicating the absence of any intrinsic magnetic moment in the unperturbed 1L-SiC lattice. This spin degeneracy is expected for a non-magnetic semiconductor and serves as a baseline for understanding the spin-polarized states that emerge when defects are introduced. The band dispersion characteristics show remarkable qualitative similarity across all functionals, with some quantitative differences. The valence bands exhibit strong dispersion near the K-point, particularly evident in the topmost valence band. This strong dispersion indicates high hole mobility in the K-point region, suggesting that hole transport could play a significant role in the electronic properties of p-type or ambipolar 1L-SiC devices. In contrast, the conduction bands show noticeably less dispersion, especially near the CBM at the M-point. This flatter conduction band suggests lower electron mobility compared to hole mobility, an asymmetry that could have important implications for charge transport in 1L-SiC-based devices. The lower electron mobility might result in longer electron lifetimes, which could be advantageous for applications requiring sustained electron excitation, such as in photocatalysis or certain optoelectronic devices. While the overall band structure features are consistent across functionals, there are significant quantitative differences, particularly in the predicted bandgap values. Table 1 provides a comprehensive summary of key electronic properties calculated using each functional, offering valuable insights into these differences. The PBE functional, employing the GGA, predicts the smallest bandgap of 2.54 eV in good agreement with [56, 54]. This underestimation is a well-known limitation of GGA functionals, stemming from their approximate treatment of exchange-correlation effects and the lack of proper self-interaction correction. Moving to more sophisticated functionals, we observe a systematic increase in the predicted bandgap. The SCAN functional, a meta-GGA approach, yields a larger bandgap of 2.81 eV, while its refined version, r2SCAN, predicts a slightly higher value of 2.86 eV. This trend reflects the improved treatment of exchange-correlation effects in meta-GGA functionals, which incorporate the kinetic energy density to better account for electron localization effects. The HSE06 hybrid functional, which includes a fraction of exact exchange, predicts the largest bandgap of 3.39 eV. Beyond the bandgap, Table 1 reveals interesting trends in other electronic properties. The work function, a critical parameter for understanding electron emission and interface properties, shows less variation across functionals compared to the bandgap. PBE predicts a work function of 4.85 eV, while SCAN and r2SCAN predict slightly higher values of 4.99 eV and 5.03 eV, respectively. Interestingly, HSE06 predicts a slightly lower work function of 4.79 eV. The electron affinity, which measures the energy required to add an electron to the neutral system, shows a non-monotonic trend across functionals. PBE and HSE06 predict similar values (2.46 eV and 2.47 eV, respectively), while SCAN and r2SCAN predict lower values (2.36 eV and 2.31 eV). This suggests that meta-GGA functionals predict a lower tendency for 1L-SiC to accept electrons compared to both PBE and HSE06. The ionization energy follows a different pattern, with SCAN and r2SCAN predicting the highest values (5.29 eV and 5.28 eV), while PBE and HSE06 predict lower values (5.10 eV and 5.03 eV). These differences in electron affinity and ionization energy could significantly influence the material’s chemical reactivity, charge transfer processes, and behavior in electronic devices. The introduction of silicon vacancies in 1L-SiC induces profound changes in its electronic structure, as illustrated in Figure 3 and supported by DOS analysis in Figure S1. For the neutral vacancy V (Figure 3a), PBE predicts a metallic state with spin-degenerate bands crossing the Fermi level, corresponding to a broad, unsplit DOS peak at the Fermi energy. These bands, derived primarily from dangling bond states of neighboring carbon atoms, exhibit minimal dispersion ( 0.2 eV bandwidth), indicating strong spatial localization of the defect states. This localization is evidenced in the DOS by sharp peaks with full width at half maximum of approximately 0.3 eV. In contrast, SCAN and r2SCAN open a small indirect gap of 1.25 eV, with the VBM at the point and the CBM at the M point. This gap opening reflects the improved treatment of self-interaction errors in meta-GGA functionals, manifested in the DOS as a clear splitting of the defect states into occupied and unoccupied regions separated by approximately 1.25 eV. The HSE06 functional reveals a markedly different electronic structure for V. Our calculations show that the spin-up bands exhibit a gap of approximately 2.0 eV between the VBM at and the CBM at K, corresponding to well-separated DOS peaks centered at -1.0 eV and +1.0 eV relative to the Fermi level. The spin-down bands, in contrast, show metallic behavior with states crossing the Fermi level, producing a sharp DOS peak at the Fermi energy with intensity approximately twice that of the spin-up peaks. This spin polarization, absent in PBE and progressively emerging in SCAN and r2SCAN results, demonstrates the crucial role of exact exchange in capturing the correct spin physics of localized defect states. The DOS shows a clear spin splitting of 0.7 eV between majority and minority spin channels, providing a quantitative measure of the magnetic moment. For the negatively charged silicon vacancy (V), shown in Figure 3b, all functionals predict more pronounced changes in the electronic structure. PBE shows increased metallicity compared to the neutral case, with multiple band crossings at the Fermi level producing a broad DOS feature spanning approximately 1 eV around the Fermi energy. This can be interpreted as the additional electron populating previously unoccupied defect states, increasing the density of states at the Fermi level. SCAN and r2SCAN, however, predict the opening of a gap in both spin channels, with a notable spin splitting evidenced by separated DOS peaks at -0.8 eV and +0.4 eV relative to the Fermi level. The HSE06 results for V exhibit the most dramatic spin polarization. A large indirect gap of 2.2 eV appears in the spin-up channel, while the spin-down channel shows a smaller gap of 0.9 eV. The DOS reveals sharply defined peaks split by 1.3 eV between spin channels, the largest spin splitting observed among all vacancy configurations. This strong spin dependence indicates that the additional electron preferentially occupies one spin channel, maximizing the exchange energy and resulting in a high-spin configuration. The DOS peaks show minimal broadening ( 0.2 eV width), consistent with the highly localized nature of these states. The positively charged vacancy V (Figure 3c) presents yet another distinct electronic landscape. PBE predicts a return to near spin degeneracy with metallic behavior, reflected in overlapping spin-up and spin-down DOS peaks at the Fermi level. SCAN and r2SCAN maintain some degree of spin polarization, with small spin-dependent gaps producing DOS peak splittings of 0.3-0.5 eV. The HSE06 functional, however, predicts a wide gap of 3.0 eV for spin-up states and metallic behavior for spin-down states, creating a highly asymmetric DOS with a large gap in one spin channel and a sharp peak at the Fermi level in the other. This stark asymmetry between spin channels implies a complex redistribution of electrons upon positive charging, possibly involving significant lattice relaxation and rehybridization of surrounding atoms, as evidenced by the redistribution of DOS weight between s and p orbital contributions compared to the neutral case. The DOS orbital analysis reveals that silicon vacancy states are dominated by p-orbital contributions, with p:s peak intensity ratios exceeding 3:1 in the defect state region. This strong p-orbital character persists across all charge states, though the relative weights shift with charging: the p-orbital contribution increases to nearly 85% in V but decreases to approximately 70% in V, reflecting charge-state-dependent changes in the local electronic structure. Carbon vacancies demonstrate distinctive electronic characteristics that significantly differ from silicon vacancies, reflecting the fundamental differences in C-Si and Si-Si bonding in the SiC lattice. For the neutral carbon vacancy (V), HSE06 calculations predict a unique electronic structure with a spin splitting of 0.3 eV, substantially smaller than the 0.7 eV observed for V. This is evidenced by both the band structure gaps (1.5 eV for spin-up and 1.2 eV for spin-down channels) and the corresponding DOS peaks separated by 0.3 eV near the Fermi level. The defect states show broader dispersion of approximately 0.5 eV across the Brillouin zone, particularly notable along the -K direction, compared to the much narrower 0.2 eV dispersion observed for silicon vacancies. The DOS analysis reveals that carbon vacancy states have significant contributions from both s and p orbitals, with p-orbital peaks showing approximately 50% greater intensity than s-orbital peaks in the energy range of -1 to 1 eV around the Fermi level. This mixed orbital character explains the enhanced band dispersion and suggests stronger hybridization with neighboring atoms compared to the predominantly p-orbital character of silicon vacancy states. The negatively charged carbon vacancy (V) exhibits a complex electronic structure that varies significantly across functionals. PBE predicts metallic behavior with multiple band crossings at the Fermi level, while SCAN and r2SCAN reveal the emergence of small gaps (0.4-0.6 eV) in both spin channels. The HSE06 calculations demonstrate spin-split gaps of 0.8 eV and 0.5 eV for spin-up and spin-down channels respectively, notably smaller than the corresponding gaps in V (2.2 eV and 0.9 eV). The DOS shows three distinct peaks at -0.8 eV, -0.2 eV, and +0.3 eV relative to the Fermi level, corresponding to specific transitions in the band structure. For the positively charged state (V), HSE06 calculations reveal a direct gap of 0.6 eV at the K point for the spin-up channel, while maintaining metallic character in the spin-down channel. The DOS exhibits a sharp peak at +0.4 eV above the Fermi level in the spin-down channel, with an integrated intensity approximately twice that of the neutral state peaks, indicating enhanced state density available for optical transitions. The progression of electronic structure across charge states shows systematic trends in the DOS peak positions: the main defect state peaks shift from -0.3 eV in V to 0.0 eV in V to +0.4 eV in V relative to the Fermi level, demonstrating clear charge state-dependent electronic structure evolution. This detailed analysis of band structures and DOS reveals that carbon vacancies maintain moderate spin polarization (0.3-0.8 eV splittings) across charge states, but with consistently smaller magnitudes than silicon vacancies (0.7-1.3 eV splittings). The broader bandwidth of carbon vacancy states (0.5 eV vs 0.2 eV) and mixed s-p orbital character indicate fundamentally different bonding characteristics, leading to more delocalized defect states that could influence their optical and magnetic properties.
III.3 Migration of Vacancies
Figure 5 presents our investigation into vacancy migration in monolayer SiC using nudged elastic band (NEB) calculations, employing the HSE06 functional with D3 dispersion correction. Figures 5(a) and 5(b) illustrate the top view of the monolayer SiC structure, showing the migration pathways for Si and C vacancies, respectively. For both vacancy types, we considered three distinct pathways: Path A (green) involves migration to the diagonal position of the same atom type within a single hexagon, Path B (pink) represents migration to the same atom type position diagonally crossing two hexagons, and Path C (blue) depicts migration to the nearest neighbor of the same atom type. The energy profiles along these migration pathways for Si vacancies (Fig. 5c) and C vacancies (Fig. 5d) reveal significant differences in migration barriers. For Si vacancies, Path C exhibits the lowest energy barrier of 0.38 eV, making it the most favorable migration route. Path A presents an intermediate barrier of 1.12 eV, while Path B has the highest barrier of 1.41 eV, making it the least likely route for Si vacancy migration. Similarly, for C vacancies, Path C again presents the lowest energy barrier at 0.81 eV, followed by Path A with 1.33 eV and Path B with 1.76 eV. These results indicate that migration to the nearest neighbor (Path C) is energetically favored for both Si and C vacancies, while migration across two hexagons (Path B) is the least likely. Notably, the energy barriers for C vacancies are consistently higher than those for Si vacancies across all pathways, suggesting that C vacancies are less mobile and more stable in the 1L-SiC lattice. The lower energy barriers for Path C can be attributed to its shorter migration distance and the preservation of the local bonding environment. In contrast, Path B involves significant bond breaking and reforming over a longer distance, leading to higher energy barriers. Path A represents an intermediate case, requiring some bond rearrangement but over a shorter distance compared to Path B. Examination of the transition state (TS) configurations provides further insights into the atomic-scale mechanisms of vacancy migration. For both vacancy types, the TS for Path C involves a displacement of the migrating atom towards the vacancy site by approximately 0.8 , with a slight out-of-plane distortion of about 0.2 . This minimal distortion contributes to the lower energy barrier for this path. In contrast, the TS for Path B shows significant in-plane lattice distortion, with neighboring atoms displaced by up to 0.5 , and out-of-plane buckling of up to 0.4 , consistent with its higher energy barrier. Interestingly, the energy difference between the initial state (IS) and final state (FS) for each migration pathway is relatively small (less than 0.1 eV), indicating that the vacancies do not have a strong preference for specific lattice sites. This suggests that vacancy movement in monolayer SiC is primarily influenced by kinetic factors rather than thermodynamic considerations. It’s important to note that our calculations consider only neutral vacancies. The migration behavior of charged vacancies may differ significantly due to changes in the local electronic structure and lattice relaxation. While we can speculate that positively charged vacancies might exhibit lower migration barriers due to reduced electron density in the defect region, and negatively charged vacancies might have higher barriers due to increased electron-electron repulsion, these predictions are purely speculative and would require further computational validation to confirm. The comparison with bulk SiC provides valuable context for understanding the unique properties of defects in 2D SiC. In 3C-SiC, for example, Si vacancy migration barriers have been reported to be around 3.5-5.2 eV [57, 58], significantly higher than our predicted value of 0.38 eV for 1L-SiC. These theoretical predictions of vacancy migration are particularly relevant given recent experimental demonstrations of 2D SiC synthesis. Chemical exfoliation methods have shown that 2D SiC is stable under ambient conditions [30], while epitaxial growth has demonstrated stability up to 1200°C in vacuum [31]. The predicted migration barriers help explain this remarkable thermal stability, though substrate effects may further modify defect dynamics in experimental samples.
III.4 Optical Properties
The introduction of vacancies in monolayer SiC (1L-SiC) induces profound changes in its optical properties, as revealed by our calculations of the dielectric function and absorption coefficient. Figure 6 presents a comprehensive analysis of the real () and imaginary () parts of the dielectric function (Figures 6a and 6c), as well as the corresponding optical absorptivity (Figures 6b and 6d), computed using the PBE functional within DFPT formalism. While HSE06 was used for electronic structure calculations, optical properties were computed using PBE-DFPT due to its established reliability in predicting optical response functions and significantly lower computational cost for excited state properties, following the approach validated by Lan et al. [59]. For pristine 1L-SiC, exhibits a prominent peak at approximately 4 eV, corresponding to fundamental optical transitions across the bandgap, in agreement with previous theoretical predictions [60]. The real part of the dielectric function () shows a characteristic oscillator-like behavior, with its magnitude directly influencing the material’s refractive index. The relationship between these dielectric components and the absorption coefficient () can be expressed through:
, where
[…]

### v2

@@ 3.2 Electronic Properties
| HSE06 | 3.39 | 4.79 | 2.47 | 5.03 |
Moving to more sophisticated functionals, we observe a systematic increase in the predicted bandgap. The SCAN functional, a meta-GGA approach, yields a larger bandgap of 2.81 eV, while its refined version, r2SCAN, predicts a slightly higher value of 2.86 eV. This trend reflects the improved treatment of exchange-correlation effects in meta-GGA functionals, which incorporate the kinetic energy density to better account for electron localization effects. The HSE06 hybrid functional, which includes a fraction of exact exchange, predicts the largest bandgap of 3.39 eV. This value is expected to be the most accurate among the functionals considered, as hybrid functionals are known to provide more reliable bandgap predictions for semiconductors due to their improved treatment of the self-interaction error. Beyond the bandgap, Table 1 reveals interesting trends in other electronic properties. The work function, a critical parameter for understanding electron emission and interface properties, shows less variation across functionals compared to the bandgap. PBE predicts a work function of 4.85 eV, while SCAN and r2SCAN predict slightly higher values of 4.99 eV and 5.03 eV, respectively. Interestingly, HSE06 predicts a slightly lower work function of 4.79 eV compared to the meta-GGA functionals. The electron affinity follows a non-monotonic trend across functionals. PBE and HSE06 predict similar values (2.46 eV and 2.47 eV, respectively), while SCAN and r2SCAN predict lower values (2.36 eV and 2.31 eV). This suggests that meta-GGA functionals predict a lower tendency for 1L-SiC to accept electrons compared to both PBE and HSE06. The ionization energy demonstrates a different pattern, with SCAN and r2SCAN predicting the highest values (5.29 eV and 5.28 eV), while PBE and HSE06 predict lower values (5.10 eV and 5.03 eV). These differences in electron affinity and ionization energy could significantly influence predictions of the material’s chemical reactivity, charge transfer processes, and behavior in electronic devices. The introduction of silicon vacancies in 1L-SiC induces profound changes in its electronic structure, as illustrated in Figure 3 and supported by density of states (DOS) analysis in the Supplementary Information (Figure S4-S5). For the neutral vacancy V (Figure 3a), our calculations reveal significant functional dependence in the predicted electronic structure. Using the PBE functional, we observe a metallic state with bands crossing the Fermi level. These bands, derived primarily from dangling bond states of neighboring carbon atoms, exhibit minimal dispersion ( 0.2 eV bandwidth), indicating strong spatial localization of the defect states. This localization is evidenced in the DOS by sharp peaks with full width at half maximum of approximately 0.3 eV.
In contrast, SCAN and r2SCAN reveal the emergence of spin polarization, with differing band structures for spin-up and spin-down channels. Both meta-GGA functionals predict small gaps in the spin-up channel, while maintaining partially occupied states in the spin-down channel. This spin polarization, absent in PBE results, demonstrates the importance of improved exchange-correlation treatment for capturing the magnetic properties of defect states. The HSE06 functional reveals the most dramatic electronic structure, with the spin-up bands exhibiting a gap of approximately 2.0 eV between the VBM at and the defect states, while the spin-down bands show metallic character with states crossing the Fermi level. To verify convergence of these electronic states, we performed multiple calculations starting from different initial densities and with various initial magnetic moments, confirming that the HSE06 result represents the true ground state. The DOS analysis reveals a clear spin splitting of 0.7 eV between majority and minority spin channels in HSE06 calculations, providing a quantitative measure of the magnetic moment. This substantial spin splitting, progressively developing from PBE to HSE06, underscores the critical role of exact exchange in correctly describing the spin physics of localized defect states. For the negatively charged silicon vacancy (V), shown in Figure 3b, we observe even more pronounced electronic structure changes across functionals. With PBE, the additional electron leads to increased occupation of defect states without significant modification of their dispersion. However, SCAN and r2SCAN both predict the opening of distinct gaps in the spin-up and spin-down channels, with clear spin polarization evident in the band structure. The HSE06 results for V exhibit the most dramatic spin polarization. A large indirect gap of 2.2 eV appears in the spin-up channel, while the spin-down channel shows a smaller gap of 0.9 eV. This spin splitting of 1.3 eV is the largest observed among all vacancy configurations studied, indicating that the additional electron enhances the magnetic character of the defect. The DOS reveals sharply defined peaks split by 1.3 eV between spin channels, with minimal broadening ( 0.2 eV width), consistent with the highly localized nature of these states. This strong spin-dependence indicates that the additional electron preferentially occupies one spin channel, maximizing the exchange energy and resulting in a high-spin configuration. For the positively charged vacancy V (Figure 3c), we observe yet another distinct electronic structure. PBE calculations predict reduced spin polarization compared to the neutral case, with the removal of an electron leading to more symmetric occupation of spin channels. SCAN and r2SCAN maintain moderate spin polarization, with spin-dependent gaps of varying magnitude. The HSE06 functional predicts a wide gap of 3.0 eV for spin-up states and metallic behavior for spin-down states, creating a highly asymmetric electronic structure. This stark asymmetry between spin channels implies a complex redistribution of electrons upon positive charging, possibly involving significant lattice relaxation and rehybridization of surrounding atoms. Orbital analysis of the defect states reveals that silicon vacancy states are dominated by p-orbital contributions from the surrounding carbon atoms, with p:s peak intensity ratios exceeding 3:1 in the defect state region. This strong p-orbital character persists across all charge states, though the relative weights shift with charging: the p-orbital contribution increases to nearly 85% in V but decreases to approximately 70% in V, reflecting charge-state-dependent changes in the local electronic structure.
Carbon vacancies demonstrate distinctive electronic characteristics that significantly differ from silicon vacancies, reflecting the fundamental differences in C-Si and Si-Si bonding in the SiC lattice. Figure 4 presents the band structures for carbon vacancies in different charge states, calculated with our four functionals. For the neutral carbon vacancy (V), PBE calculations predict a small spin splitting, with slightly different band structures for the two spin channels. SCAN and r2SCAN enhance this spin polarization, with more pronounced differences between spin-up and spin-down bands, particularly near the Fermi level.
HSE06 calculations reveal a unique electronic structure with a moderate spin splitting of 0.3 eV, substantially smaller than the 0.7 eV observed for V. This is evidenced by both the band structure gaps (1.5 eV for spin-up and 1.2 eV for spin-down channels) and the corresponding DOS peaks separated by 0.3 eV near the Fermi level. The defect states show broader dispersion of approximately 0.5 eV across the Brillouin zone, particularly notable along the -K direction, compared to the much narrower 0.2 eV dispersion observed for silicon vacancies. This broader bandwidth indicates stronger hybridization with neighboring atoms and a more delocalized character of the defect states. DOS analysis reveals that carbon vacancy states have significant contributions from both s and p orbitals of the surrounding silicon atoms, with p-orbital peaks showing approximately 50% greater intensity than s-orbital peaks in the energy range of -1 to 1 eV around the Fermi level. This mixed orbital character explains the enhanced band dispersion and suggests stronger hybridization with neighboring atoms compared to the predominantly p-orbital character of silicon vacancy states. The negatively charged carbon vacancy (V) exhibits a complex electronic structure that varies significantly across functionals. PBE predicts metallic behavior with multiple band crossings at the Fermi level, while SCAN and r2SCAN reveal the emergence of small gaps (0.4-0.6 eV) in both spin channels.
The HSE06 calculations demonstrate spin-split gaps of 0.8 eV and 0.5 eV for spin-up and spin-down channels respectively, notably smaller than the corresponding gaps in V (2.2 eV and 0.9 eV). The DOS shows three distinct peaks at -0.8 eV, -0.2 eV, and +0.3 eV relative to the Fermi level, corresponding to specific transitions in the band structure. For the positively charged state (V), HSE06 calculations reveal a direct gap of 0.6 eV at the K point for the spin-up channel, while maintaining metallic character in the spin-down channel. The DOS exhibits a sharp peak at +0.4 eV above the Fermi level in the spin-down channel, with an integrated intensity approximately twice that of the neutral state peaks, indicating enhanced state density available for optical transitions. The progression of electronic structure across charge states shows systematic trends in the DOS peak positions: the main defect state peaks shift from -0.3 eV in V to 0.0 eV in V to +0.4 eV in V relative to the Fermi level, demonstrating clear charge state-dependent electronic structure evolution. This detailed analysis of band structures and DOS reveals that carbon vacancies maintain moderate spin polarization (0.3-0.8 eV splittings) across charge states, but with consistently smaller magnitudes than silicon vacancies (0.7-1.3 eV splittings). The broader bandwidth of carbon vacancy states (0.5 eV vs 0.2 eV) and mixed s-p orbital character indicate fundamentally different bonding characteristics, leading to more delocalized defect states that could influence their optical and magnetic properties.
3.3 Migration of Vacancies
Figure 5 presents our comprehensive investigation of vacancy migration mechanisms in monolayer SiC using nudged elastic band (NEB) calculations with the HSE06 functional including D3 dispersion correction. Figures 5(a) and 5(b) illustrate the top view of the monolayer SiC structure, showing the distinct migration pathways for Si and C vacancies, respectively. For both vacancy types, we considered three mechanistically different pathways: Path A (green) represents migration to the diagonal position of the same atom type within a single hexagon, Path B (pink) follows migration to the same atom type position diagonally crossing two hexagons, and Path C (blue) describes migration to the nearest neighbor of the same atom type. The energy profiles along these migration pathways for Si vacancies (Fig. 5c) and C vacancies (Fig. 5d) reveal significant differences in migration barriers. For Si vacancies, our calculations show that Path C exhibits the lowest energy barrier of 0.8 eV, making it the most favorable migration route. Path A presents an intermediate barrier of 1.25 eV, while Path B shows the highest barrier of 1.75 eV, making it the least likely route for Si vacancy migration. These values represent significant revisions to previously reported barriers, reflecting our more accurate treatment of transition state geometries and careful convergence of the NEB calculations. For C vacancies, the calculated migration barriers are: Path C presents the lowest energy barrier at approximately 1.0 eV, followed by Path A with approximately 1.2 eV, and Path B with approximately 1.4 eV. The consistently higher barriers for C vacancies compared to Si vacancies suggest that C vacancies are fundamentally less mobile in the 1L-SiC lattice. Examination of the transition state (TS) configurations provides crucial insights into the atomic-scale mechanisms of vacancy migration. For Si vacancies following Path C, the TS involves a displacement of the migrating silicon atom towards the vacancy site by approximately 0.8 , accompanied by a modest out-of-plane distortion of 0.2 . The small magnitude of these distortions contributes to the relatively low barrier height. The transition state exhibits C2v symmetry, maintaining two mirror planes perpendicular to the SiC sheet. In contrast to Path C, the transition state for Path B shows substantially larger structural distortions. For Si vacancies, the TS structure exhibits significant in-plane lattice distortion, with neighboring atoms displaced by up to 0.5 , and more pronounced out-of-plane buckling reaching 0.4 . This extensive reorganization explains the higher barrier of 1.75 eV. The transition state loses all mirror symmetry elements, adopting a C1 point group that reflects the complex atomic rearrangements required for migration across two hexagons. The surrounding atoms show collective distortions extending to second and third nearest neighbors, indicating the long-range nature of the structural response. Path A presents an intermediate case, with the transition state showing moderate distortions. For Si vacancies, the migrating atom undergoes a displacement of approximately 0.6 in-plane and 0.3 out-of-plane, maintaining a single mirror plane (Cs symmetry). The resulting barrier of 1.25 eV reflects this intermediate level of structural reorganization. Notably, the transition state retains partial symmetry, unlike the lowest-symmetry configuration found in Path B. For carbon vacancies, the transition states show systematically larger distortions across all pathways, correlating with their higher migration barriers. The Path C transition state for C vacancies involves a 1.0 displacement of the migrating carbon atom, with out-of-plane distortion reaching 0.3 . This increased distortion compared to Si vacancies (0.8 and 0.2 respectively) helps explain the barrier of approximately 1.0 eV. The stronger covalent bonding character of carbon atoms requires more energy for bond breaking and reformation during migration. The Path B transition state for C vacancies shows the most extreme structural changes, with in-plane distortions reaching 0.7 and out-of-plane buckling of 0.6 . This extensive reconstruction results in the barrier of approximately 1.4 eV. The large distortions reflect the significant energy cost of breaking multiple C-Si bonds simultaneously during migration across two hexagons. The transition state geometry shows complex rehybridization of carbon orbitals, evidenced by significant changes in C-Si bond angles from the ideal 120o to values ranging from 105o to 135o. A comparative analysis of Si and C vacancy migration mechanisms reveals fundamental differences in their behavior. Silicon vacancies generally maintain more rigid local structures during migration, with smaller distortions of the surrounding lattice. This rigidity can be attributed to the more ionic character of Si-C bonds and the larger size of the silicon atom, which results in smaller relative displacements during migration. The energy cost for breaking and reforming Si-C bonds is also lower, contributing to the generally smaller barriers for Si vacancy migration.
Carbon vacancies, by contrast, induce more pronounced structural distortions during migration due to the stronger covalent character of their bonds and the smaller atomic radius of carbon. The migration process for carbon vacancies involves significant rehybridization of orbitals and more substantial rearrangement of the surrounding lattice, resulting in consistently higher energy barriers. These differences in migration barriers have important implications for defect dynamics in 1L-SiC. he significantly lower barrier for Si vacancy migration (0.8 eV via Path C) compared to C vacancies (approximately 1.0 eV via Path C) suggests that Si vacancies will be more mobile at moderate temperatures. Using transition state theory with a typical attempt frequency of 1013 s-1, we can estimate diffusion rates at various temperatures. At room temperature (300 K), Si vacancies (with 0.8 eV) would have a hopping rate of approximately 2.5 × 10-1 s-1. If the barrier for C vacancies (Path C) is approximately 1.0 eV, their hopping rate would be around 1.5 × 10-8 s-1. At higher temperatures such as 600 K, the difference becomes less pronounced but still significant: Si vacancies would have a hopping rate on the order of 105- 106 s-1, while C vacancies would diffuse at about 3.6 × 104 s-1. These differences in mobility could lead to preferential aggregation of Si vacancies or formation of vacancy complexes under thermal annealing conditions, potentially influencing the evolution of defect structures in 2D SiC. The higher mobility of Si vacancies may also explain the experimental observation of divacancy complexes (V-V) in SiC,14 as mobile Si vacancies could migrate towards relatively immobile C vacancies and form stable complexes.
3.4 Optical Properties
The introduction of vacancies in monolayer SiC (1L-SiC) induces profound changes in its optical properties. Figure 6 presents our analysis of the dielectric response and absorption characteristics, computed using density functional perturbation theory (DFPT) within the PBE framework. We calculated both the real () and imaginary () parts of the dielectric function (Figures 6a and 6c), along with the corresponding optical absorptivity (Figures 6b and 6d). For pristine 1L-SiC, shows a prominent peak at approximately 4 eV, corresponding to fundamental optical transitions across the bandgap. The real part of the dielectric function () exhibits characteristic oscillator behavior, with its magnitude directly influencing the material’s refractive index. For the neutral silicon vacancy (V), we observe a distinct peak at 0.1 eV in both and , with reaching approximately 2-3. This peak position aligns with transitions between the defect states identified in our electronic structure calculations. The absorption coefficient of V shows enhancement of approximately 3-4% around 0.07 eV. The negatively charged silicon vacancy (V) produces a broad response below 0.1 eV in both and , with multiple peaks of magnitude around 5. This electronic structure results in enhanced absorption in the far-infrared region, showing peaks of 3-4% at energies below 0.1 eV. These features correlate directly with the multiple transitions available between spin-split states identified in our electronic structure analysis. For the positively charged state (V), we find the strongest dielectric response, with showing peaks reaching 20 at energies below 0.1 eV. This translates to the highest absorption among all vacancy states, with enhancement up to 12% in the far-infrared region. The absorption profile shows multiple distinct bands between 0.1-0.3 eV, reflecting transitions between the asymmetric spin channels observed in the electronic structure. These multiple absorption features correspond to transitions between the well-defined defect levels in the band structure of V shown in Figure 3c. Carbon vacancies demonstrate distinctly different optical responses compared to silicon vacancies, reflecting their more delocalized electronic structure. For the neutral carbon vacancy (V), we observe moderate peaks in ( 4), resulting in absorption coefficients of 3-4%. These relatively weak oscillator strengths correlate with the broader dispersion of carbon vacancy states identified in our band structure analysis. The absorption features appear more distributed across the far-infrared region, lacking the sharp peaks characteristic of silicon vacancies. The negatively charged carbon vacancy (V) shows enhanced absorption (5-6%) compared to the neutral state, arising from additional transitions enabled by the extra electron. The dielectric response exhibits three distinct features corresponding to the DOS peaks identified at -0.8 eV, -0.2 eV, and +0.3 eV relative to the Fermi level. These transitions produce a more complex absorption spectrum than V, with multiple overlapping bands below 0.2 eV, reflecting the more delocalized nature of the defect states. The positively charged state (V) demonstrates the strongest response among carbon vacancy configurations, with reaching approximately 20 below 0.05 eV. This enhancement results in absorption peaks exceeding 20%, making it particularly promising for far-infrared applications. Unlike the discrete absorption bands seen in V, V exhibits a more continuous absorption profile extending up to 0.4 eV, reflecting its more delocalized electronic structure.
[…]